I found it quite tough..
Log_x 9=7+Log_3 x^2+2Log_3 x

Answer by stanbon(75887)

Log_x 9=7+Log_3 x^2+2Log_3 x
---
logx(9) = 7 + log3(x^2)+log3(x)
-----
logx(9) = 7 + log3[x^3]
---
Using a calculator graph
Y1 = log(9)/log(x)
and
Y2 = 7 + [log(x^3)/log(3)]
---
Then see where they intersect.
I get:
x = 1.326721..
Cheers,
Stan H.
===================